# EvAlignUX: Advancing UX Research through LLM-Supported Exploration of Evaluation Metrics

### v1

@@ 6. Discussion
In this section, we explore methods for promoting a mindset-focused UX evaluation in HCI education and practice, examine the potential risks of using large language models for evaluation, and discuss potential design implications yielded from our findings.
6.1. From Method-Centric to Mindset-Centric: a Paradigm Shift of Creating UX Evaluation Plans
@@
In this study, we found that highlighting UX metrics as a center of UX evaluation empowers UX researchers holistically, enhancing both their perceived quality of their proposed plans and their confidence in their UX evaluation plans (RQ1). Furthermore, the combined use of different EvAlignUX components for refining their UX plans allowed participants to self-validate, critique, and rationalize their evaluation choices (RQ2), with their perceived inspirations from the system varying according to their UX experience (RQ3). Meanwhile, we observed their thought processes evolving from asking general questions to using AI to support iterative reasoning for their chosen metrics, considering various types of information, including previously not considered aspects, such as outcomes and risks.
UX evaluation plays a vital role in ensuring the quality of HCI system research, and it is a growing field with diverse approaches (Pettersson et al., 2018). Although several UX evaluation methods have been outlined in previous literature reviews (Pettersson et al., 2018; Vermeeren et al., 2010; Rivero and Conte, 2017; Darin et al., 2019; Kieffer et al., 2019), few interaction designs have been proposed to support UX researchers in developing comprehensive UX evaluation plans. Gray’s previous work (Gray, 2016) found that many UX designers and practitioners view UX as a set of rigid methods rather than a ”mindset,” suggesting that HCI education should move beyond teaching predefined methods and instead focus on fostering the adaptability and improvisation skills necessary for UX designers to address the unique challenges they face. A mindset approach to UX evaluation, as emphasized in our findings, aligns with their argument.
Several scholars emphasize adaptability and context-sensitive practices in UX research. Vermeeren et al. (Vermeeren et al., 2010) note that no single UX evaluation method suits all projects, urging researchers to adapt methods to specific contexts. Rivero et al. (Rivero and Conte, 2017) advocate for iterative frameworks that incorporate feedback, while Darin et al. (Darin et al., 2019) stress the importance of critical reflection throughout the evaluation process. Building on these insights, our study highlights the role of metrics in establishing a mindset-oriented approach to UX evaluation. This shift is essential for advancing the field and ensuring UX evaluations meaningfully contribute to the success of HCI systems research.
6.2. Design Implications: Using LLM for UX for Interdisciplinary and Crossed-Studies Research
Our findings reveal that EvAlignUX provides five key benefits for developing UX research plans. However, we observed varying usage patterns among participants (RQ1, RQ2): some enriched system-generated plans with detailed context, others focused on applying specific metrics, and some simply copied the plan without critical evaluation. We found that these patterns were, to some extent, linked to participants’ UX experience (RQ3).
@@ 6.2. Design Implications: Using LLM for UX for Interdisciplinary and Crossed-Studies Research
While some participants viewed EvAlignUX as a potential “teammate” in developing UX plans, the challenge remains in ensuring that LLM usage does not lead to the deskilling of UX practitioners’ research expertise. Previous literature has highlighted the ethical concerns and potential harms of over-reliance on LLMs in various domains. For instance, Ahmad (Ahmad et al., 2023b) found that excessive dependence on AI in education led to students becoming passive, diminishing their decision-making abilities and negatively impacting cognitive skills. Similarly, in human-AI co-creative writing, AI outputs are heavily influenced by the distribution of their training datasets (Roselli et al., 2019; Argyle et al., 2023). Without considering users’ unique needs, over-reliance on these outputs can result in research homogeneity and a lack of creativity (Lee et al., 2024a; Rezwana and Maher, 2023). Additionally, issues such as hallucination, lack of transparency, and algorithmic bias have been identified as major barriers to effective AI collaboration (Zhai et al., 2024). Given the limitations of LLMs, future studies need to carefully consider and address the ethical concerns raised to support the critical thinking of UX researchers. Here are some potential research directions:
@@ 6.2.1. Mitigating Potential Over-Reliance and Triggering Deeper Thought Processes
One potential direction is to use AI systems that provide hints, or to use Question Bank (see Fig. 8), this approach may aim to augment researchers’ reasoning by prompting reflection and encouraging deeper engagement with their work. Essential features include controllability, and explainability, alongside clear disclaimers that warn users about the risks of over-reliance on generative AI (Liao et al., 2020b). Previous research highlights that transparent systems with clear AI outputs foster better human-AI collaboration by encouraging users to engage more critically with the tool (Wang et al., 2020; Glinka and Müller-Birn, 2023). Furthermore, studies indicate that embedding mechanisms that encourage active thinking—such as presenting alternative viewpoints or highlighting uncertainties—can help mitigate the risk of over-dependence on AI-generated solutions and foster deeper cognitive engagement (Bansal et al., 2021).
@@ 6.2.2. Using AI to Support Self-Critique
Our findings suggest that AI may assist in addressing challenges through self-critique mechanisms. For example, EvAlignUX uses its literature base to provide users with relevant information that encourages self-critique. However, unlike human experts, AI systems often fail to identify nuanced flaws in research plans that arise from researchers’ misunderstandings. Previous studies have shown that AI’s capacity for self-critique is limited, particularly when dealing with complex, contextual errors that require domain-specific knowledge or critical thinking (Bansal et al., 2021; Wang et al., 2021). AI’s ability to critique existing research, while promising, lacks the contextual depth found in human-to-human collaboration (Wang et al., 2021; Xu and Mulligan, 2022). Further research is needed to explore how AI can more closely mimic human critical evaluation in the research process.
@@ 6.2.3. Using AI to Mitigate UX Risks
A key design implication often overlooked is the concept of “impact visioning.” While only one participant in our study considered the potential risks during the initial development of their UX plans, previous studies have shown that negative consequences of new technologies can often be predicted in early design stages (Lee et al., 2024b; Pang et al., 2024). AI systems, by drawing on cross-disciplinary insights, may also help researchers anticipate and mitigate potential risks in UX evaluation, raising awareness of emerging concerns in technology use and design. Prior studies in AI ethics and risk management suggest that AI can be instrumental in identifying unintended consequences and risks in various domains, including UX design, by leveraging data across disciplines to foresee issues that might not be immediately apparent (Vinuesa et al., 2020; Floridi and Chiriatti, 2021).
@@ 6.3. Limitations and Future Work
We acknowledge several limitations that may affect both EvAlignUX ’s performance and the findings of our research. First, as an exploratory system, EvAlignUX incorporates a limited selection of papers, primarily focused on human-AI interaction. Participants proposing studies in interdisciplinary fields may find the system’s suggestions less relevant or accurate. Second, although we observed differences in participants’ perceptions and usage behaviors, the majority of participants were doctoral students, which may limit the generalizability of our findings. Third, for the EvAlignUX backend, the recommendation is confined to metrics that have been predefined in our database. Future studies should explore the continued expansion of the database to include a broader spectrum of metrics and online publication sources.
@@ 7. Conclusion
User experience is vital to any Human-Computer Interaction research, yet many UX designers and practitioners often rely on rigid evaluation methods rather than adopting a holistic approach. In this paper, we introduced EvAlignUX , an LLM-based tool designed to support UX evaluation development. Through an evaluation with 19 HCI researchers, we found that centering UX metrics in the evaluation process fostered a mindset that balanced holistic UX impact with contextual factors, evaluation practices, and research value. This approach significantly improved participants’ perceived plan quality and confidence, while also prompting their thought processes to consider risks when evaluating systems. For practical implications, we offer a Question Bank to guide future UX evaluation efforts; and suggest that future researchers can enhance EvAlignUX by incorporating domain-specific literature to further refine their evaluation planning. In a world where experience defines impact, a UX-driven mindset serves as the key to meaningful and lasting designs.
@@ Acknowledgements.
This material is based upon work supported by the National Science Foundation under Grant No. 2119589. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Additionally, results presented in this paper were obtained using CloudBank (Norman et al., 2021), which is supported by the National Science Foundation under award No. 1925001.References
@@
- Fan et al. (2022) Mingming Fan, Xianyou Yang, TszTung Yu, Q Vera Liao, and Jian Zhao. 2022. Human-ai collaboration for UX evaluation: effects of explanation and synchronization. Proceedings of the ACM on Human-Computer Interaction 6, CSCW1 (2022), 1–32.
@@
- Norman et al. (2021) Michael Norman, Vince Kellen, Shava Smallen, Brian DeMeulle, Shawn Strande, Ed Lazowska, Naomi Alterman, Rob Fatland, Sarah Stone, Amanda Tan, et al. 2021. CloudBank: Managed Services to Simplify Cloud Access for Computer Science Research and Education. In Practice and Experience in Advanced Research Computing. 1–4.
@@
- Rezwana and Maher (2023) Jeba Rezwana and Mary Lou Maher. 2023. User Perspectives on Ethical Challenges in Human-AI Co-Creativity: A Design Fiction Study. In Proceedings of the 15th Conference on Creativity and Cognition (Virtual Event, USA) (C&C ’23). Association for Computing Machinery, New York, NY, USA, 62–74. https://doi.org/10.1145/3591196.3593364
@@
- Wang et al. (2020) Dakuo Wang, Elizabeth Churchill, Pattie Maes, Xiangmin Fan, Ben Shneiderman, Yuanchun Shi, and Qianying Wang. 2020. From human-human collaboration to Human-AI collaboration: Designing AI systems that can work together with people. In Extended abstracts of the 2020 CHI conference on human factors in computing systems. 1–6.

### v2

@@ 6. Discussion
In this section, we discuss methods for fostering a mindset-focused UX evaluation in HCI education and practice, the potential risks of using Large Language Models to support UX evaluation, and the design implications of our findings.
6.1. Shifting UX Evaluation from Method-Centric to Mindset-Centric
Gray’s prior work (Gray, 2016) highlighted that many UX practitioners view UX as a ”rigid set of methods” rather than as a ”mindset”. This method-centric perspective often emphasizes the application of techniques without fully aligning with project-specific goals. Similarly, prior research (Goodman, 2013; McLellan, 2000; Roedl and Stolterman, 2013; Gray et al., 2014) has critiqued academic discussions for focusing on method creation and testing, neglecting real-world implementation contexts. This disconnect creates a ”projected practice community” (Gray et al., 2014; Gray, 2016), where academic conceptions fail to align with practitioners’ realities.
@@ 6.1. Shifting UX Evaluation from Method-Centric to Mindset-Centric
The interaction design and evaluation of EvAlignUX demonstrated that it empowered UX researchers to conduct UX evaluation more holistically. First, providing metric definitions, relevant literature, and examples that enabled for researchers to reflect on the implications of these metrics enhanced both the perceived quality of their proposed plans and their confidence in their UX evaluation plans (RQ1). Second, combining different EvAlignUX components to refine UX plans allowed participants to self-validate, critique, and rationalize their evaluation choices (RQ2). Third, participants’ thought processes and perceived inspirations from the system varied based on their UX experience and evolved from asking general questions to leveraging AI for iterative reasoning about their chosen metrics, including previously unconsidered aspects, such as outcomes and AI risks (RQ3).
6.1.1. Mindset-Centric Evaluation: Encouraging Reflections on Insights
Unlike method-centric approaches, mindset-centric evaluation emphasizes critical reflection on why, what, and how evaluations are conducted, aligning choices with broader implications–goals, values, and the insights researchers and practitioners aim to achieve. This paradigm shift acknowledges the complexity and adaptability required in UX evaluation, bridging the gap between academic research and practical applications. EvAlignUX facilitates this shift by not only helping define metrics, enabling users to understand their underlying concepts, relevance, and alignment with other research areas, but also by connecting researchers with example research papers and real-world applications. This integration encouraged them to move beyond focusing solely on specific methods and to instead reflect on the insights they aim to achieve, such as fostering engagement or building trustworthiness. By shifting their question from ”What method should I use?” to ”What insights am I seeking?”, EvAlignUX helped researchers adopt a more holistic perspective. Additionally, the inclusion of AI risk considerations encouraged users to evaluate their UX plans through new lenses, adding depth and broadening the implications of their work. This shift enriched the researchers’ mental models, underscoring EvAlignUX ’s value in promoting a mindset-centric approach to UX evaluation.
6.1.2. Implications for HCI Practice and Education
This mindset-centric shift holds significant implications for both HCI practice and education.
First, it aligns with the increasing diversity and sophistication of UX evaluation approaches (Pettersson et al., 2018). While numerous reviews outline UX evaluation methods (Pettersson et al., 2018; Vermeeren et al., 2010; Rivero and Conte, 2017; Darin et al., 2019; Kieffer et al., 2019), few interaction designs directly support the development of comprehensive evaluation plans. For example, scholars have stressed the need for adaptable, context-sensitive approaches: Vermeeren et al. (Vermeeren et al., 2010) stress that no single method fits all projects; Rivero et al. (Rivero and Conte, 2017) advocate for iterative frameworks that incorporate feedback; and Darin et al. (Darin et al., 2019) call for continuous critical reflection throughout the evaluation process. Our design aligns with the intention to democratize the design process: when EvAlignUX users select metrics from the recommended options, click “more” to explore additional metrics, edit the index to update the recommendations, or adjust their selection after identifying potential risks, they provide feedback to the system. In the final UX evaluation export, users can further refine their selected metrics. Our findings demonstrate an effective interaction design that fosters a mindset-centric perspective, informing and elevating HCI systems research.
Second, our findings reveal that EvAlignUX offers five key benefits for developing UX research plans. However, participants exhibited diverse usage patterns (RQ1, RQ2). Some enriched system-generated plans with detailed context; others prioritized applying the application of specific metrics; and some adopted the plans without critical evaluation. These patterns were, to some extent, linked to participants’ UX experience (RQ3). For example, less experienced UX researchers used EvAlignUX to explore unfamiliar metrics, domains, and terminologies, thus finding it more inspirational. In contrast, experienced researchers primarily leveraged the system as a validation tool rather than a source of inspiration. This suggests that EvAlignUX ’s primary value evolves with users’ expertise—serving as an exploratory and instructive tool for novices while functioning as a confirming and refining aid for experts. Prior work emphasizes that HCI education should extend beyond teaching predefined methods to fostering adaptability, improvisation, and critical thinking skills (Gray, 2016). EvAlignUX embodies these principles by enabling researchers to tailor their approaches to the unique demands of each project. Rather than enforcing rigid processes, future AI-driven UX evaluation systems and educational frameworks should encourage researchers to treat metrics as core design materials—fundamental elements that shape and guide evaluation. By adopting this mindset-centric perspective, researchers and practitioners can better navigate evolving complexities, ensuring their UX evaluations remain meaningful, impactful, and contextually relevant.
Third, the literature-based UX repository is a growing collection that could expand to include industry case studies and metrics, broadening its applicability beyond academic research. While academic UX research emphasizes systematic methodologies, theoretical frameworks, and empirical validation, industry UX case studies often prioritize practical insights, real-world constraints, and agile decision-making processes. Integrating industry contributions would make the repository more reflective of diverse UX practices, enabling designers, engineers, and customers to generate tailored plans and share them for collaborative decision-making. Future studies could explore how to structure this integration effectively, curating a broader spectrum of metrics from both peer-reviewed sources and real-world implementations, while maintaining clarity and usability. Additionally, incorporating industry case studies could introduce new ways of evaluating UX beyond traditional academic benchmarks, offering a more holistic perspective on how different sectors define and measure UX.
6.2. Towards Responsible Use of LLMs in Interdisciplinary and Cross-Disciplinary UX Research
@@
Previous literature has highlighted the ethical concerns and potential harms of over-reliance on LLMs in various domains. For instance, Ahmad et al. (2023b) found that excessive dependence on AI in education led to students becoming passive, diminishing their decision-making abilities, and negatively impacting their cognitive skills. Similarly, in human-AI co-creative writing, AI outputs are heavily influenced by the distribution of their training datasets (Roselli et al., 2019; Argyle et al., 2023). Without considering users’ unique needs, over-reliance on these outputs can result in research homogeneity and a lack of creativity (Lee et al., 2024a; Rezwana and Maher, 2023). Additionally, issues such as hallucination, lack of transparency, and algorithmic bias have been identified as major barriers to effective AI collaboration (Zhai et al., 2024). To mitigate these risks, future iterations of EvAlignUX and similar tools could be designed with adaptive modes or adjustable feedback mechanisms that align with a user’s experience level. Such tailoring would help novices develop core competencies over time while allowing more experienced users to refine advanced strategies, ultimately supporting meaningful skill growth rather than hindering it. Given the limitations of LLMs, future studies need to carefully consider and address the ethical concerns raised, ensuring that UX researchers remain critical, reflective, and ultimately self-reliant evaluators. Potential research directions include:
@@ 6.2.1. Mitigating Potential Over-Reliance and Triggering Deeper Thought Processes
Questions about potential over-reliance have also emerged. We position the design of such a tool as an educational resource that potentially bridges formal education and real-world practice, helping novices transition into more advanced roles while also supporting experts in adopting new perspectives. It is designed to foster a mindset-centric approach, which inherently encourages reflection and critical thinking, counteracting de-skilling risks. One potential direction is to use AI systems that provide hints, or to use a Question Bank (see Fig. 9). This approach may aim to augment researchers’ reasoning by prompting reflection and encouraging deeper engagement with their work. Essential features include controllability and explainability, alongside clear disclaimers that warn users about the risks of over-reliance on generative AI (Liao et al., 2020b). Previous research highlights that transparent systems with clear AI outputs foster better human-AI collaboration by encouraging users to engage more critically with the tool (Wang et al., 2020a; Glinka and Müller-Birn, 2023). Furthermore, studies indicate that embedding mechanisms that encourage active thinking—such as presenting alternative viewpoints or highlighting uncertainties—can help mitigate the risk of over-dependence on AI-generated solutions and foster deeper cognitive engagement (Bansal et al., 2021).
@@ 6.2.2. Using AI to Support Self-Critique
Our findings suggest that AI may assist in addressing challenges through self-critique mechanisms. For example, EvAlignUX uses its literature base to provide users with relevant information that encourages self-critique. However, unlike human experts, AI systems often fail to identify nuanced flaws in research plans that stem from researchers’ misunderstandings. Previous studies have shown that AI’s capacity for self-critique is limited, particularly when dealing with complex, contextual errors that require domain-specific knowledge or critical thinking (Bansal et al., 2021; Wang et al., 2021). While AI’s ability to critique existing research is promising, it lacks the contextual depth found in human-to-human collaboration (Wang et al., 2021; Xu and Mulligan, 2022). Further research is needed to explore how AI can more closely mimic human critical evaluation in the research process.
@@ 6.2.3. Using AI to Mitigate UX Risks
A key design implication often overlooked is the concept of “impact visioning.” While only one participant in our study considered potential risks during the initial development of their UX plans, previous studies have shown that the negative consequences of new technologies can often be predicted in early design stages (Lee et al., 2024b; Pang et al., 2024). AI systems, by drawing on cross-disciplinary insights, may also help researchers anticipate and mitigate potential risks in UX evaluation, raising awareness of emerging concerns in technology use and design. Prior studies in AI ethics and risk management suggest that AI can be instrumental in identifying unintended consequences and risks in various domains, including UX design, by leveraging data across disciplines to foresee issues that might not be immediately apparent (Vinuesa et al., 2020; Floridi and Chiriatti, 2021). As the body of evaluation metrics and AI-related risks expands in the rapidly evolving landscape of human-AI interaction, practitioners can benefit from recommendations that continuously evolve.
@@ 6.3. Limitations and Future Work
We acknowledge several limitations that can affect EvAlignUX ’s performance and the findings. First, as an exploratory system, it incorporates a limited selection of papers, primarily focused on human-AI interaction. Also, for the EvAlignUX backend, recommendations are confined to metrics that have been predefined in our database. Participants proposing studies in interdisciplinary fields may find the system’s suggestions less relevant or accurate; however, we do not claim that the current metrics repository is exhaustive. Second, although we observed differences in participants’ perceptions and usage behaviors, the majority of participants were doctoral students, which may limit the generalizability of our findings. Our user evaluation was conducted by HCI researchers, but in the future, we can include industry practitioners who may lack formal UX research training but need to conduct effective system evaluations. Expanding the system’s design to better support these groups in conducting effective evaluations could be an important direction for future development. Finally, EvAlignUX scaffolds the conceptual phase of early-stage UX research planning. It focuses on ideation, metric selection, and critical reflection, providing a foundation for subsequent stages, such as methods, logistics, and participant considerations. For instance, selecting metrics upfront facilitates targeted discussions about participants and data collection methods. However, we acknowledge that detailed logistical support is beyond the tool’s scope and could be addressed in future enhancements.
@@ 7. Conclusion
User experience is vital to any Human-Computer Interaction research, yet many UX designers and practitioners often rely on rigid evaluation methods rather than adopting a holistic approach. In this paper, we introduced EvAlignUX , an LLM-based tool designed to support UX evaluation development. Through an evaluation with 19 HCI researchers, we found that centering UX metrics in the evaluation process fostered a mindset that balanced holistic UX impact with contextual factors, evaluation practices, and research value. This approach significantly improved participants’ perceived plan quality and confidence, while also prompting their thought processes to consider risks when evaluating systems. For practical implications, we offer a Question Bank to guide future UX evaluation efforts, and suggest that future researchers can enhance EvAlignUX by incorporating domain-specific literature to further refine their evaluation planning. In a world where experience defines impact, a UX-driven mindset serves as the key to meaningful and lasting designs evaluation.
@@ Acknowledgements.
This material is based upon work supported by the National Science Foundation under Grant #2119589. The work is also supported under the AI Research Institutes program by the National Science Foundation and the Institute of Education Sciences, U.S. Department of Education through Award #2229873 - AI Institute for Transforming Education for Children with Speech and Language Processing Challenges. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation, the Institute of Education Sciences, or the U.S. Department of Education. Additionally, this project is made possible in part by the Institute of Museum and Library Services RE-252329-OLS-22. The views, findings, conclusions or recommendations expressed in this article do not necessarily represent those of the Institute of Museum and Library Services. The authors acknowledge the National Artificial Intelligence Research Resource (NAIRR) Pilot (NAIRR240177) for contributing to this research. We sincerely thank the anonymous reviewers for their constructive suggestions, which have significantly strengthened this work.References
@@
- Goodman (2013) Elizabeth Sarah Goodman. 2013. Delivering design: Performance and materiality in professional interaction design. University of California, Berkeley.
- Graesser et al. (2006) Arthur C Graesser, Zhiqiang Cai, Max M Louwerse, and Frances Daniel. 2006. Question Understanding Aid (QUAID) a web facility that tests question comprehensibility. Public Opinion Quarterly 70, 1 (2006), 3–22.
@@
- Gray et al. (2014) Colin M Gray, Erik Stolterman, and Martin A Siegel. 2014. Reprioritizing the relationship between HCI research and practice: bubble-up and trickle-down effects. In Proceedings of the 2014 conference on Designing interactive systems. 725–734.
@@
- Hassenzahl et al. (2010) Marc Hassenzahl, Sarah Diefenbach, and Anja Göritz. 2010. Needs, affect, and interactive products–Facets of user experience. Interacting with computers 22, 5 (2010), 353–362.
@@
- Hsiao et al. (2014) I-Han Hsiao, Shuguang Han, Manav Malhotra, Hui Soo Chae, and Gary Natriello. 2014. Survey sidekick: Structuring scientifically sound surveys. In Intelligent Tutoring Systems: 12th International Conference, ITS 2014, Honolulu, HI, USA, June 5-9, 2014. Proceedings 12. Springer, 516–522.
@@
- Kim et al. (2019) Soomin Kim, Joonhwan Lee, and Gahgene Gweon. 2019. Comparing data from chatbot and web surveys: Effects of platform and conversational style on survey response quality. In Proceedings of the 2019 CHI conference on human factors in computing systems. 1–12.
@@
- McLellan (2000) Hilary McLellan. 2000. Experience design. Cyberpsychology and behavior 3, 1 (2000), 59–69.
@@
- Nielsen (1994a) Jakob Nielsen. 1994a. Enhancing the explanatory power of usability heuristics. In Proceedings of the SIGCHI conference on Human Factors in Computing Systems. 152–158.
- Nielsen (1994b) Jakob Nielsen. 1994b. Usability engineering. Morgan Kaufmann.
@@
- Norman (2007) Don Norman. 2007. Emotional design: Why we love (or hate) everyday things. Basic books.
@@
- Rezwana and Maher (2023) Jeba Rezwana and Mary Lou Maher. 2023. User Perspectives on Ethical Challenges in Human-AI Co-Creativity: A Design Fiction Study. In Proceedings of the 15th Conference on Creativity and Cognition (Virtual Event, USA) (Creativity and Cognition ’23). Association for Computing Machinery, New York, NY, USA, 62–74. https://doi.org/10.1145/3591196.3593364
@@
- Roedl and Stolterman (2013) David J Roedl and Erik Stolterman. 2013. Design research at CHI and its applicability to design practice. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. 1951–1954.
@@
- Siemon et al. (2016) Dominik Siemon, Taras Rarog, and Susanne Robra-Bissantz. 2016. Semi-automated questions as a cognitive stimulus in idea generation. In 2016 49th Hawaii International Conference on System Sciences (HICSS). IEEE, 257–266.
@@
- Wan et al. (2024) Hongyu Wan, Jinda Zhang, Abdulaziz Arif Suria, Bingsheng Yao, Dakuo Wang, Yvonne Coady, and Mirjana Prpa. 2024. Building LLM-based AI Agents in Social Virtual Reality. In Extended Abstracts of the CHI Conference on Human Factors in Computing Systems. 1–7.
- Wang et al. (2020a) Dakuo Wang, Elizabeth Churchill, Pattie Maes, Xiangmin Fan, Ben Shneiderman, Yuanchun Shi, and Qianying Wang. 2020a. From human-human collaboration to Human-AI collaboration: Designing AI systems that can work together with people. In Extended abstracts of the 2020 CHI conference on human factors in computing systems. 1–6.
@@
- Wang et al. (2020b) Jinping Wang, Hyun Yang, Ruosi Shao, Saeed Abdullah, and S Shyam Sundar. 2020b. Alexa as coach: Leveraging smart speakers to build social agents that reduce public speaking anxiety. In Proceedings of the 2020 CHI conference on human factors in computing systems. 1–13.
@@
- Willis and Lessler (2013) GB Willis and JT Lessler. 2013. Question appraisal system: QAS 99. National Cancer Institute.
- Willis (2004) Gordon B Willis. 2004. Cognitive interviewing: A tool for improving questionnaire design. sage publications.